# How Far Are We on the Decision-Making of LLMs? Evaluating LLMs’ Gaming Ability in Multi-Agent Environments

### v3

@@ 5.1 Specific Games
Other than papers listed in Table 4 on evaluating LLMs using classical games, researchers have explored diverse scenarios involving more complicated games. Using the complex and deceptive environments of Avalon game as a test bed, recent work focuses on long-horizon multi-party dialogues Stepputtis et al. (2023), social behaviors Lan et al. (2023), and recursive contemplation Wang et al. (2023) for identifying deceptive information. Other papers have investigated the application of LLMs in communication games like Werewolf, with a focus on tuning-free frameworks Xu et al. (2023b) and reinforcement learning-powered approaches Xu et al. (2023c). O’Gara (2023) found that advanced LLMs exhibit deception and lie detection capabilities in the text-based game, Hoodwinked. Meanwhile, Liang et al. (2023a) evaluated LLMs’ intelligence and strategic communication skills in the word guessing game, Who Is Spy? In the game of Water Allocation Challenge, Mao et al. (2023) constructed a scenario highlighting unequal competition for limited resources.
@@ 5.2 Game Benchmarks
Another line of studies collects games to build more comprehensive benchmarks to assess the artificial general intelligence of LLMs. Tsai et al. (2023) found that while LLMs, such as ChatGPT, perform competitively in text games, they struggle with world modeling and goal inference. GameEval Qiao et al. (2023) introduced three goal-driven conversational games (Ask-Guess, SpyFall, and TofuKingdom) to effectively assess the problem-solving capabilities of LLMs in cooperative and adversarial settings. MAgIC Xu et al. (2023a) proposed the probabilistic graphical modeling method for evaluating LLMs in multi-agent game settings. LLM-Co Agashe et al. (2023) developed the LLM-Coordination framework to assess LLMs in multi-agent coordination scenarios, showcasing their capabilities in partner intention inference and proactive assistance. Wu et al. introduced SmartPlay, a benchmark for evaluating LLMs as agents across six games, emphasizing reasoning, planning, and learning capabilities. These papers focus on games with more complex designs, while our study investigates eight classical and essential games in game theory.
@@ Appendix D Examples of GPT-4-Rephrased Prompts
§4.1 involves testing the GPT-3.5 (0125)’s robustness against different prompt templates. This section presents the prompts used in this analysis, namely Prompts V2 to V4, with V1 as the default, as detailed in §C). We include only the prompts for the game “Guess 2/3 of the Average,” while the five prompt templates of seven other games can be found in the supplementary materials.
@@ Appendix E Rescale Method for Raw Scores
$\begin{split}S_{1}&=\begin{cases}\frac{MAX-S_{1}}{MAX-MIN}*100,&R<1\\ \frac{\lvert 2S_{1}-(MAX-MIN)\rvert}{MAX-MIN}*100,&R=1\\ \frac{S_{1}}{MAX-MIN}*100,&R>1\end{cases},\\ S_{2}&=\frac{\max(R,1-R)-S_{2}}{\max(R,1-R)}*100,\\ S_{3}&=\frac{G-S_{3}}{G}*100,\\ S_{4}&=\frac{T-S_{4}}{T}*100,\\ S_{5}&=(1-S_{5})*100,\\ S_{6}&=\frac{S_{6}}{\max_{ij}v_{ij}}*100,\\ S_{7}&=S_{7}*100,\\ S_{8}&=\frac{2*G-S_{8P}}{2*G}*50+S_{8V}*50.\\ \end{split}$ (1)
@@ F.1 Robustness: Multiple Runs
Public Goods Game $41.3$ $25.4$ $45.7$ $38$ $44.0$ $38.8_{\pm 8.1}$
@@
Sealed-Bid Auction $6.5$ $6.5$ $4.6$ $5.5$ $4.4$ $5.5_{\pm 1.0}$
Battle Royale $20.0$ $21.4$ $46.7$ $23.5$ $31.3$ $28.6_{\pm 11.0}$
Pirate Game $80.5$ $71.0$ $72.0$ $74.8$ $59.8$ $71.6_{\pm 7.6}$
Overall $44.9$ $40.7$ $46.2$ $42.5$ $43.2$ $43.5_{\pm 2.2}$
@@ F.2 Robustness: Temperatures
Divide the Dollar $69.3$ $67.0$ $67.7$ $67.9$ $72.8$ $68.1$ $68.8_{\pm 2.1}$
Public Goods Game $15.3$ $10.8$ $17.9$ $18.0$ $36.5$ $41.3$ $23.3_{\pm 12.5}$
@@
Pirate Game $75.0$ $54.0$ $77.8$ $84.0$ $59.8$ $80.5$ $71.8_{\pm 12.1}$
Overall $37.2$ $35.8$ $40.9$ $39.1$ $42.4$ $44.9$ $40.4_{\pm 3.4}$
@@
F.3 Robustness: Prompt Versions
@@
Divide the Dollar $68.1$ $81.0$ $91.5$ $75.8$ $79.7$ $79.2_{\pm 8.5}$
Public Goods Game $41.3$ $26.6$ $45.2$ $50.2$ $24.2$ $37.5_{\pm 11.5}$
@@ F.4 Reasoning Strategies
Public Goods Game $41.3$ $56.1$ $32.4$ $37.4$ $25.6$
Diner’s Dilemma $31.0$ $82.5$ $0.0$ $17.5$ $47.0$
Sealed-Bid Auction $6.5$ $1.4$ $7.4$ $4.0$ $5.8$
Battle Royale $20.0$ $17.6$ $6.3$ $33.3$ $26.7$
Pirate Game $80.5$ $71.0$ $80.5$ $74.8$ $59.8$
Overall $44.9$ $57.4$ $42.5$ $36.8$ $45.7$
@@ F.6 Leaderboard
F.7 Detailed Player Actions of GPT-3.5 (0125)
@@
Appendix H Limitations
@@
This study is subject to several limitations. Firstly, due to time and budget constraints, we do not evaluate all prominent LLMs such as LLaMA, Vicuna, Mixtral, and Falcon. However, we promise to expand our leaderboard to include more LLMs in the future. Secondly, our experiments do not explore scenarios where different LLMs compete in the same game. Instead, our evaluation uses ten agents derived from the same LLM. We acknowledge that including diverse LLMs in the same game could yield more intriguing insights. This aspect is designated for a future direction. Thirdly, we limit the games to 20 rounds and inform the agents of this total, potentially affecting strategies in Betraying games where agents may collaborate initially and betray in the final round for greater gain. We also leave this part as our future research agenda.
Appendix I Ethics Statement and Broader Impacts
Currently, our paper excludes human interactions and user studies. We use no existing assets, thereby avoiding copyright issues. Our research seeks to evaluate and enhance LLMs’ reasoning capabilities, promoting their application in decision-making scenarios. Our framework potentially benefits society by enabling humans to engage with LLMs in gameplay. Enhanced reasoning capabilities in LLMs could position them as decision-making assistants for humans.

### v5

@@ 5 Related Work
Evaluating LLMs through game theory models has become a popular research direction. An overview on recent studies is summarized in Table 3. We find: (1) Many studies examine the PSNE on two-player, single-round settings, focusing on the Prisoner’s Dilemma and the Ultimatum Game. (2) Varying temperatures are employed without discussing the impact on LLMs’ performance.
@@ 5.1 Specific Games
Researchers have explored diverse game scenarios. Using the complex and deceptive environments of Avalon game as a test bed, recent work focuses on long-horizon multi-party dialogues Stepputtis et al. (2023), social behaviors Lan et al. (2023), social intelligence Liu et al. (2024), and recursive contemplation Wang et al. (2023) for identifying deceptive information. Other papers have investigated communication games like Werewolf, with a focus on tuning-free frameworks Xu et al. (2023b) and reinforcement learning-powered approaches Xu et al. (2024). O’Gara (2023) found that advanced LLMs exhibit deception and lie detection capabilities in the text-based game, Hoodwinked. Meanwhile, Liang et al. (2023) evaluated LLMs’ intelligence and strategic communication skills in the word guessing game, Who Is Spy? In the game of Water Allocation Challenge, Mao et al. (2023) constructed a scenario highlighting unequal competition for limited resources.
@@ 5.2 Game Benchmarks
Another line of studies collects games to build more comprehensive benchmarks to assess the artificial general intelligence of LLMs. Tsai et al. (2023) found that while LLMs perform competitively in text games, they struggle with world modeling and goal inference. GameEval Qiao et al. (2023) introduced three goal-driven conversational games (Ask-Guess, SpyFall, and TofuKingdom) to assess the problem-solving capabilities of LLMs in cooperative and adversarial settings. MAgIC Xu et al. (2023a) proposed the probabilistic graphical modeling method for evaluating LLMs in multi-agent game settings. LLM-Co Agashe et al. (2023) assesses LLMs in multi-agent coordination scenarios, showcasing their capabilities in partner intention inference and proactive assistance. SmartPlay Wu et al. (2024) evaluated LLMs as agents across six games, emphasizing reasoning, planning, and learning capabilities. Abdelnabi et al. (2024) designed negotiation games involving six parties with distinct objectives to evaluate LLMs’ ability to reach agreement.
@@ 6 Conclusion
Acknowledgments
The paper is supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14206921 of the General Research Fund).
@@ Appendix D Examples of GPT-4-Rephrased Prompts
§4.1 involves testing the GPT-3.5 (0125)’s robustness against different prompt templates. This section presents the prompts used in this analysis, namely Prompts V2 to V4, with V1 as the default, as detailed in §C). We include only the prompts for the game “Guess 2/3 of the Average,” while the five prompt templates of seven other games can be found in our GitHub (https://github.com/CUHK-ARISE/GAMABench).
@@ Appendix E Rescale Method for Raw Scores
$\begin{split}S_{1}&=\begin{cases}\frac{(MAX-MIN)-S_{1}}{MAX-MIN}*100,&R<1\\ \left(1-\frac{\lvert 2S_{1}-(MAX-MIN)\rvert}{MAX-MIN}\right)*100,&R=1\\ \frac{S_{1}}{MAX-MIN}*100,&R>1\end{cases},\\ S_{2}&=\frac{\max(R,1-R)-S_{2}}{\max(R,1-R)}*100,\\ S_{3}&=\max\left(\frac{G-S_{3}}{G}*100,0\right),\\ S_{4}&=\begin{cases}\frac{T-S_{4}}{T}*100,&\frac{R}{N}\leq 1\\ \frac{S_{4}}{T}*100,&\frac{R}{N}>1\end{cases},\\ S_{5}&=(1-S_{5})*100,\\ S_{6}&=S_{6}*100,\\ S_{7}&=S_{7}*100,\\ S_{8}&=\frac{2*G-S_{8P}}{2*G}*50+S_{8V}*50.\\ \end{split}$ (1)
@@ F.1 Robustness: Multiple Runs
Public Goods Game $41.2$ $25.4$ $45.7$ $38.0$ $44.0$ $38.9_{\pm 8.1}$
@@
Sealed-Bid Auction $14.6$ $14.6$ $11.6$ $12.9$ $11.5$ $13.0_{\pm 1.5}$
Battle Royale $20.0$ $21.4$ $46.7$ $23.5$ $31.2$ $28.6_{\pm 11.0}$
Pirate Game $80.6$ $71.2$ $72.0$ $74.7$ $59.5$ $71.6_{\pm 7.7}$
Overall $45.9$ $41.7$ $47.1$ $43.4$ $44.1$ $44.4_{\pm 2.1}$
@@ F.2 Robustness: Temperatures
Divide the Dollar $69.3$ $67.0$ $67.6$ $67.9$ $72.8$ $68.1$ $68.8_{\pm 2.1}$
Public Goods Game $15.3$ $10.7$ $17.8$ $18.0$ $36.5$ $41.2$ $23.3_{\pm 12.5}$
@@
Pirate Game $75.0$ $53.9$ $77.7$ $83.8$ $59.5$ $80.6$ $71.7_{\pm 12.1}$
Overall $38.1$ $36.7$ $41.9$ $39.9$ $43.2$ $45.9$ $41.0_{\pm 3.4}$
@@
F.3 Robustness: Prompt Templates
@@
Divide the Dollar $68.1$ $81.0$ $91.4$ $75.8$ $79.6$ $79.2_{\pm 8.5}$
Public Goods Game $41.2$ $26.6$ $45.2$ $50.2$ $24.2$ $37.5_{\pm 11.5}$
@@
Public Goods Game $41.2$ $56.1$ $32.4$ $37.4$ $25.6$
Diner’s Dilemma $4.0$ $82.5$ $0.0$ $17.5$ $47.0$
Sealed-Bid Auction $14.6$ $5.3$ $16.3$ $11.6$ $13.0$
Battle Royale $20.0$ $17.6$ $6.2$ $33.3$ $26.7$
Pirate Game $80.6$ $71.2$ $80.6$ $74.7$ $59.5$
Overall 45.9 57.9 43.7 37.8 46.5
@@ F.5 Leaderboard
F.6 Detailed Player Actions of GPT-3.5 (0125)
@@
The above experiments implicitly assume that players are not informed about others’ fixed strategies. To investigate the effect of explicit information, we design additional experiments using the “Guess 2/3 of the Average” game, where players are provided varying levels of information about others’ strategies:
  1. Setting (a).
    The player is explicitly informed that others are smart and will always choose 0 (the Nash equilibrium).
  2. Setting (b).
    The player is informed that others are smart but not explicitly told they will always choose the Nash equilibrium.
  3. Setting (c).
    The player is informed that others are stupid and will choose random numbers.
These experiments are conducted using GPT-4o, and the results are as follows:
  1. Setting (a).
    GPT-4o selects 0 in the first round and continues to do so in all subsequent rounds.
  2. Setting (b).
    GPT-4o does not select 0 in the first round but converges to 0 within a few rounds.
  3. Setting (c).
    GPT-4o’s selections are random, indicating an inability to infer the optimal choice of $33$ (calculated as $50\times\frac{2}{3}$, given the average of others’ random selections is $50$).
Appendix H Jailbreak Influence
To bypass the value alignment in LLMs, we use the jailbreak technique, specifically CipherChat Yuan et al. (2024). Prior research demonstrated that this method can exacerbate negative traits in GPT-4 Huang et al. (2024b). To assess whether value alignment influences the behavior of LLMs, we evaluate GPT-4o’s performance in behavioral contexts before and after applying CipherChat, using the Public Goods Game and the Diner’s Dilemma.
Prior to CipherChat, GPT-4o achieves scores of $90.91\pm 2.72$ in the Public Goods Game and $10.7\pm 8.3$ in the Diner’s Dilemma. After the jailbreak, its performance declines to $88.55\pm 2.38$ and $7.5\pm 4$1, respectively. This decline reflects GPT-4o’s inherent risk-averse tuning. For instance, in the Public Goods Game, GPT-4o prioritizes minimizing losses, reasoning, “If no contribution is made to the public pot, I will have no loss.” Similarly, in the Diner’s Dilemma, it opts for the less costly dish to reduce expenditures. A comparable conservative approach is observed in the El Farol Bar Game, where GPT-4o tends to avoid the risk of overcrowding by staying home. In conclusion, GPT-4o adopts riskier strategies after jailbreak, such as contributing less in the Public Goods Game and selecting the more expensive dish in the Diner’s Dilemma.
One assumption is that CipherChat could reduce the model’s prosocial behavior, make it more self-serving, and increase its scores in our betraying games. We believe that the observed decrease in scores results from OpenAI’s efforts to enhance GPT-4o’s value alignment, thereby mitigating the influence of CipherChat. Following Huang et al. (2024b), we assess the model’s negative traits using the Dark Triad Dirty Dozen. The results of GPT-4o, both before and after the jailbreak, are presented in Table 9. Contrary to the findings of Huang et al. (2024b), which reported increased scores after jailbreak, our analysis indicates a decrease in these negative traits for GPT-4o. This result suggests that GPT-4o does not exhibit heightened negative characteristics, such as selfishness, even after being subjected to jailbreak attempts.
GPT-4o w/o Jailbreak w/ Jailbreak
Machiavellianism $4.6\pm 0.4$ $3.5\pm 1.3$
Psychopathy $3.5\pm 0.4$ $3.0\pm 0.5$
Neuroticism $6.1\pm 0.4$ $5.2\pm 0.8$
Table 9: The jailbroken GPT-4o’s results on Dark Triad Dirty Dozen.
Appendix I Limitations
@@
This study is subject to several limitations. Firstly, due to time and budget constraints, we do not evaluate all prominent LLMs such as LLaMA-3.2, Qwen-2.5 and Claude-3.5. However, we promise to expand our leaderboard to include more LLMs in the future. Secondly, our experiments do not explore scenarios where different LLMs compete in the same game. Instead, our evaluation uses ten agents derived from the same LLM. We acknowledge that including diverse LLMs in the same game could yield more intriguing insights. This aspect is designated for a future direction. Thirdly, we limit the games to 20 rounds and inform the agents of this total, potentially affecting strategies in Betraying games where agents may collaborate initially and betray in the final round for greater gain. We also leave this part as our future research agenda. However, we believe 20 rounds are sufficient to observe agent behavior patterns. Extending the rounds exceeds the token limit without yielding new observations, as the convergence trend remains consistent.
Appendix J Ethics Statement and Broader Impacts
Our research seeks to evaluate and enhance LLMs’ reasoning capabilities, facilitating their application in decision-making scenarios. On the one hand, users need to notice that current LLMs often display self-interested behavior in decision-making, which may not maximize social welfare. On the other hand, our framework promotes societal benefits by facilitating human-LLM interaction through gameplay, which can be applied in educational contexts such as economics and game theory. Ultimately, enhancing LLMs’ reasoning skills could enable them to serve as effective decision-making assistants for humans.
Appendix K LLaMA-3.1-70B
K.1 Robustness: Multiple Runs
Tests T1 (Default) T2 T3 T4 T5 $Avg_{\pm Std}$
Guess 2/3 of the Average $82.2$ $82.7$ $84.3$ $84.6$ $86.4$ $84.0_{\pm 1.7}$
El Farol Bar $64.2$ $55.8$ $61.7$ $60.0$ $56.7$ $59.7_{\pm 3.5}$
Divide the Dollar $87.9$ $92.0$ $86.0$ $80.8$ $88.6$ $87.0_{\pm 4.1}$
Public Goods Game $93.4$ $90.8$ $84.7$ $90.4$ $93.6$ $90.6_{\pm 3.6}$
Diner’s Dilemma $47.0$ $41.5$ $56.0$ $44.5$ $51.5$ $48.1_{\pm 5.7}$
Sealed-Bid Auction $15.6$ $20.2$ $13.8$ $13.6$ $15.4$ $15.7_{\pm 2.7}$
Battle Royale $70.0$ $90.0$ $92.9$ $100.0$ $35.7$ $77.7_{\pm 26.0}$
Pirate Game $42.8$ $53.8$ $71.4$ $81.8$ $70.3$ $64.0_{\pm 15.5}$
Overall $62.9$ $65.8$ $68.8$ $69.5$ $62.3$ $65.9_{\pm 3.3}$
K.2 Robustness: Temperatures
Table 11: Quantitative results of playing the games with temperature ranging from 00 to $1$.
Temperatures 0.0 0.2 0.4 0.6 0.8 1.0 (Default) $Avg_{\pm Std}$
Guess 2/3 of the Average $75.7$ $84.7$ $80.6$ $84.9$ $83.9$ $82.2$ $82.0_{\pm 3.5}$
El Farol Bar $6.7$ $50.0$ $46.7$ $53.3$ $63.3$ $64.2$ $47.4_{\pm 21.2}$
Divide the Dollar $95.0$ $87.6$ $90.0$ $90.4$ $91.1$ $87.9$ $90.3_{\pm 2.7}$
Public Goods Game $33.8$ $79.8$ $70.8$ $83.6$ $83.0$ $93.4$ $74.0_{\pm 21.0}$
Diner’s Dilemma $28.0$ $27.0$ $34.0$ $36.5$ $45.0$ $47.0$ $36.2_{\pm 8.4}$
Sealed-Bid Auction $12.5$ $13.7$ $18.8$ $15.0$ $12.7$ $15.6$ $14.7_{\pm 2.4}$
Battle Royale $94.4$ $86.7$ $56.2$ $95.0$ $80.0$ $70.0$ $80.4_{\pm 15.1}$
Pirate Game $46.0$ $46.0$ $70.4$ $75.5$ $79.1$ $42.8$ $60.0_{\pm 16.7}$
Overall $49.0$ $59.4$ $58.4$ $66.8$ $67.3$ $62.9$ $60.6_{\pm 6.8}$
K.3 Robustness: Prompt Templates
Prompt Versions V1 (Default) V2 V3 V4 V5 $Avg_{\pm Std}$
Guess 2/3 of the Average $82.2$ $87.5$ $83.2$ $90.5$ $82.4$ $85.2_{\pm 3.7}$
El Farol Bar $64.2$ $63.3$ $63.3$ $58.3$ $64.2$ $62.7_{\pm 2.5}$
Divide the Dollar $87.9$ $95.1$ $84.1$ $87.6$ $94.0$ $89.7_{\pm 4.6}$
Public Goods Game $93.4$ $92.9$ $87.4$ $67.6$ $89.0$ $86.1_{\pm 10.6}$
Diner’s Dilemma $47.0$ $47.5$ $34.0$ $53.0$ $47.0$ $45.7_{\pm 7.0}$
Sealed-Bid Auction $15.6$ $5.4$ $13.0$ $6.1$ $10.6$ $10.2_{\pm 4.4}$
Battle Royale $70.0$ $90.0$ $75.0$ $41.2$ $85.0$ $72.2_{\pm 19.1}$
Pirate Game $42.8$ $77.0$ $88.8$ $58.6$ $73.0$ $68.1_{\pm 17.8}$
K.4 Generalizability
Guess 2/3 of the Average $Avg_{\pm Std}$
[…]
$94.1$ $91.4$ $92.0$ $83.8$ $82.2$ $81.4$ $72.6$ $89.6$ $93.0$ $92.4$ $90.3$ $89.9$ $90.9$ $88.0_{\pm 6.2}$
El Farol Bar $Avg_{\pm Std}$
[…]
$73.0$ $81.2$ $70.0$ $64.2$ $63.7$ $72.0$ $70.7_{\pm 6.5}$
Divide the Dollar $Avg_{\pm Std}$
[…]
$72.1$ $87.9$ $91.6$ $95.6$ $97.5$ $88.9_{\pm 10.1}$
Public Goods Game $Avg_{\pm Std}$
[…]
$95.4$ $95.5$ $95.3$ $93.4$ $82.9$ $92.5_{\pm 4.9}$
Diner’s Dilemma $Avg_{\pm Std}$
[…]
$47.0$ $48.5$ $44.5$ $37.5$ $31.0$ $40.0$ $41.4_{\pm 6.6}$
Sealed-Bid Auction $Avg_{\pm Std}$
[…]
$4.1$ $4.4$ $7.6$ $13.6$ $7.4_{\pm 3.8}$
Battle Royale $Avg_{\pm Std}$
[…]
$41.2$ $70.0$ $70.0$ $60.39_{\pm 13.59}$
Pirate Game $Avg_{\pm Std}$
[…]
$71.1$ $70.2$ $42.8$ $48$ $58.1_{\pm 14.7}$
Appendix L Gemini-1.5-Pro
L.1 Robustness: Multiple Runs
Tests T1 (Default) T2 T3 T4 T5 $Avg_{\pm Std}$
Guess 2/3 of the Average $96.2$ $95.4$ $95.1$ $95.1$ $95.1$ $95.4_{\pm 0.5}$
El Farol Bar $37.5$ $40.0$ $35.8$ $30.8$ $41.7$ $37.2_{\pm 4.2}$
Divide the Dollar $93.8$ $94.2$ $94.2$ $93.5$ $93.5$ $93.8_{\pm 0.3}$
Public Goods Game $100.0$ $100.0$ $100.0$ $100.0$ $100.0$ $100.0_{\pm 0.0}$
Diner’s Dilemma $29.0$ $43.0$ $33.0$ $38.5$ $36.0$ $35.9_{\pm 5.3}$
Sealed-Bid Auction $42.5$ $25.3$ $21.4$ $27.0$ $18.2$ $26.9_{\pm 9.4}$
Battle Royale $75.0$ $90.0$ $71.4$ $85.0$ $85.0$ $81.3_{\pm 7.7}$
Pirate Game $92.2$ $83.9$ $88.8$ $94.0$ $80.6$ $87.9_{\pm 5.6}$
Overall $70.8$ $71.5$ $67.5$ $70.5$ $68.8$ $69.8_{\pm 1.6}$
L.2 Robustness: Temperatures
Table 15: Quantitative results of playing the games with temperature ranging from 00 to $1$.
Temperature 0.0 0.2 0.4 0.6 0.8 1.0 $Avg_{\pm Std}$
Guess 2/3 of the Average $96.1$ $99.2$ $96.6$ $96.6$ $96.4$ $96.2$ $96.9_{\pm 1.2}$
El Farol Bar $37.5$ $20.0$ $28.3$ $40.0$ $38.3$ $37.5$ $33.6_{\pm 7.8}$
Divide the Dollar $94.5$ $93.5$ $93.5$ $94.5$ $93.3$ $93.8$ $93.8_{\pm 0.5}$
Public Goods Game $100.0$ $100.0$ $100.0$ $100.0$ $100.0$ $100.0$ $100.0_{\pm 0.0}$
Diner’s Dilemma $33.5$ $45.0$ $43.0$ $36.5$ $42.0$ $29.0$ $38.2_{\pm 6.2}$
Sealed-Bid Auction $31.1$ $24.1$ $27.9$ $21.0$ $32.4$ $42.5$ $29.8_{\pm 7.5}$
Battle Royale $88.9$ $85.0$ $80.0$ $75.0$ $87.5$ $75.0$ $81.9_{\pm 6.1}$
Pirate Game $96.0$ $90.3$ $96.1$ $99.2$ $96.0$ $92.2$ $95.0_{\pm 3.2}$
Overall $72.2$ $69.6$ $70.7$ $70.3$ $73.2$ $70.8$ $71.1_{\pm 1.3}$
L.3 Robustness: Prompt Templates
Version V1 (Default) V2 V3 V4 V5 $Avg_{\pm Std}$
Guess 2/3 of the Average $96.2$ $95.1$ $92.7$ $97.2$ $88.9$ $94.0_{\pm 3.3}$
El Farol Bar $37.5$ $53.3$ $60.8$ $46.7$ $27.5$ $45.2_{\pm 13.1}$
Divide the Dollar $93.8$ $90.3$ $62.1$ $100.0$ $92.5$ $87.7_{\pm 14.8}$
Public Goods Game $100.0$ $97.2$ $98.7$ $100.0$ $99.8$ $99.1_{\pm 1.2}$
Diner’s Dilemma $29.0$ $24.0$ $22.0$ $18.0$ $23.0$ $23.2_{\pm 4.0}$
Sealed-Bid Auction $42.5$ $38.6$ $33.5$ $8.2$ $20.5$ $28.7_{\pm 14.2}$
Battle Royale $75.0$ $92.3$ $70.0$ $75.0$ $85.0$ $79.5_{\pm 9.0}$
Pirate Game $92.2$ $82.3$ $92.3$ $82.3$ $77.8$ $85.4_{\pm 6.5}$
L.4 Generalizability
Guess 2/3 of the Average $Avg_{\pm Std}$
[…]
$98.5$ $99.4$ $98.6$ $97.8$ $95.4$ $91.1$ $5.3$ $97.0$ $97.7$ $97.3$ $92.5$ $88.0$ $75.8$ $87.3_{\pm 25.4}$
El Farol Bar $Avg_{\pm Std}$
[…]
$80.5$ $56.9$ $32.5$ $42.5$ $41.9$ $66.5$ $53.5_{\pm 17.9}$
Divide the Dollar $Avg_{\pm Std}$
[…]
$96.5$ $93.8$ $98.4$ $93.8$ $100.0$ $96.5_{\pm 2.8}$
Public Goods Game $Avg_{\pm Std}$
[…]
$100.0$ $100.0$ $100.0$ $100.0$ $100.0$ $100.0_{\pm 0.0}$
Diner’s Dilemma $Avg_{\pm Std}$
[…]
$29.0$ $12.0$ $24.5$ $11.5$ $16.5$ $42.5$ $22.7_{\pm 11.9}$
Sealed-Bid Auction $Avg_{\pm Std}$
[…]
$24.0$ $42.5$ $38.4$ $44.9$ $37.4_{\pm 9.4}$
Battle Royale $Avg_{\pm Std}$
[…]
$92.3$ $75.0$ $75.0$ $80.8_{\pm 8.2}$
Pirate Game $Avg_{\pm Std}$
[…]
$79.2$ $85.3$ $92.2$ $98.6$ $88.8_{\pm 8.4}$
Appendix M GPT-4o
M.1 Robustness: Multiple Runs
Tests T1 (Default) T2 T3 T4 T5 $Avg_{\pm Std}$
Guess 2/3 of the Average $94.9$ $94.8$ $94.2$ $94.1$ $93.4$ $94.3_{\pm 0.6}$
El Farol Bar $95.0$ $41.7$ $70.8$ $55.0$ $87.5$ $70.0_{\pm 22.1}$
Divide the Dollar $95.7$ $95.7$ $94.9$ $94.0$ $95.4$ $95.2_{\pm 0.7}$
Public Goods Game $94.1$ $88.1$ $87.4$ $93.5$ $91.5$ $90.9_{\pm 3.0}$
Diner’s Dilemma $23.5$ $4.5$ $3.5$ $8.0$ $14.0$ $10.7_{\pm 8.3}$
Sealed-Bid Auction $19.2$ $18.8$ $17.7$ $25.3$ $23.0$ $20.8_{\pm 3.2}$
Battle Royale $89.5$ $60.0$ $50.0$ $72.2$ $65.0$ $67.3_{\pm 14.8}$
Pirate Game $77.3$ $88.4$ $93.7$ $79.8$ $82.8$ $84.4_{\pm 6.7}$
Overall $73.6$ $61.5$ $64.0$ $65.2$ $69.1$ $66.7_{\pm 4.7}$
M.2 Robustness: Temperatures
Table 19: Quantitative results of playing the games with temperature ranging from 00 to $1$.
Temperature 0.0 0.2 0.4 0.6 0.8 1.0 $Avg_{\pm Std}$
Guess 2/3 of the Average $94.4$ $94.4$ $94.4$ $94.4$ $93.2$ $94.9$ $94.3_{\pm 0.6}$
El Farol Bar $66.7$ $50.8$ $44.2$ $65.0$ $75.0$ $95.0$ $66.1_{\pm 18.1}$
Divide the Dollar $100.0$ $99.1$ $98.6$ $94.3$ $97.7$ $95.7$ $97.6_{\pm 2.2}$
Public Goods Game $87.6$ $87.0$ $87.2$ $87.8$ $92.1$ $94.1$ $89.3_{\pm 3.0}$
Diner’s Dilemma $27.0$ $12.5$ $8.0$ $49.5$ $64.5$ $23.5$ $30.8_{\pm 21.9}$
Sealed-Bid Auction $24.6$ $22.6$ $24.0$ $21.2$ $22.8$ $19.2$ $22.4_{\pm 2.0}$
Battle Royale $73.7$ $50.0$ $50.0$ $20.0$ $77.8$ $89.5$ $60.2_{\pm 25.2}$
Pirate Game $99.5$ $92.7$ $88.4$ $75.8$ $82.3$ $77.3$ $86.0_{\pm 9.2}$
Overall $71.7$ $63.6$ $61.9$ $63.5$ $75.7$ $73.6$ $68.3_{\pm 6.0}$
M.3 Robustness: Prompt Templates
Version V1 (Default) V2 V3 V4 V5 $Avg_{\pm Std}$
Guess 2/3 of the Average $94.9$ $93.0$ $94.7$ $94.3$ $91.6$ $93.7_{\pm 1.4}$
El Farol Bar $95.0$ $72.5$ $37.5$ $59.2$ $60.8$ $65.0_{\pm 21.0}$
Divide the Dollar $95.7$ $95.7$ $95.6$ $93.9$ $96.1$ $95.4_{\pm 0.9}$
Public Goods Game $94.1$ $96.2$ $89.4$ $88.6$ $94.0$ $92.4_{\pm 3.3}$
Diner’s Dilemma $23.5$ $50.0$ $50.0$ $33.5$ $37.5$ $38.9_{\pm 11.3}$
Sealed-Bid Auction $19.2$ $38.1$ $35.0$ $20.3$ $33.6$ $29.2_{\pm 8.8}$
Battle Royale $89.5$ $60.0$ $10.0$ $64.7$ $30.0$ $50.8_{\pm 31.1}$
Pirate Game $77.3$ $93.7$ $67.9$ $88.9$ $86.5$ $82.9_{\pm 10.3}$
M.4 Generalizability
Guess 2/3 of the Average $Avg_{\pm Std}$
[…]
$99.3$ $98.0$ $96.6$ $95.0$ $94.9$ $88.8$ $22.7$ $55.4$ $46.2$ $72.8$ $69.1$ $76.8$ $75.0$ $76.2_{\pm 23.4}$
El Farol Bar $Avg_{\pm Std}$
[…]
$99.0$ $91.2$ $87.5$ $95.0$ $56.9$ $83.5$ $85.5_{\pm 15.1}$
Divide the Dollar $Avg_{\pm Std}$
[…]
$92.5$ $95.7$ $97.3$ $97.5$ $98.3$ $96.3_{\pm 2.3}$
Public Goods Game $Avg_{\pm Std}$
[…]
$100.0$ $95.3$ $94.4$ $88.6$ $89.8$ $93.6_{\pm 4.6}$
Diner’s Dilemma $Avg_{\pm Std}$
[…]
$23.5$ $46.0$ $10.0$ $14.5$ $2.5$ $13.0$ $18.2_{\pm 15.2}$
Sealed-Bid Auction $Avg_{\pm Std}$
[…]
$20.9$ $23.8$ $21.4$ $26.0$ $23.0_{\pm 2.3}$
Battle Royale $Avg_{\pm Std}$
[…]
$82.4$ $55.0$ $65.0$ $67.5_{\pm 13.8}$
Pirate Game $Avg_{\pm Std}$
[…]
$73.8$ $47.2$ $80.6$ $83.6$ $71.3_{\pm 16.6}$